OpenStudy (anonymous):

Evaluate \[\huge \left| \sqrt{7-4\sqrt{3}} \right|\]

OpenStudy (nincompoop):

that is interesting :) show me how you would approach the problem

OpenStudy (anonymous):

My brain thought to make that into a perfect square somehow

OpenStudy (nincompoop):

okay then try that and see

OpenStudy (anonymous):

how to do that , is there a method

OpenStudy (nincompoop):

well you said perfect square.. what is a perfect square ?

OpenStudy (anonymous):

how to make \[\huge 7-4\sqrt{3}\] into a prfect square

OpenStudy (nincompoop):

let us try a simpler problem then I will answer your question \[\sqrt{\sqrt{16}}\] solve that

OpenStudy (anonymous):

that's 4th root of 16

OpenStudy (anonymous):

2

OpenStudy (anonymous):

16^{1/4}

OpenStudy (anonymous):

=2

OpenStudy (nincompoop):

great!

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@zepdrix

OpenStudy (nincompoop):

well, even so, sometimes we need to keep going back to simpler ideas simply saying that you are in college doesn't tell me anything how deep your math knowledge is 7-4 sqrt(3) = 4-4 sqrt(3)+3 = 4-4 sqrt(3)+(sqrt(3))^2 = (2-sqrt(3))^2

OpenStudy (anonymous):

that was a smart move , thank you !

OpenStudy (nincompoop):

\[7-4 \sqrt{3} = 4-4 \sqrt{3}+3 = 4-4 \sqrt{3}+{\sqrt{3}}^2 = (2-\sqrt{3})^2\]

OpenStudy (anonymous):

that was just trial and guess right

OpenStudy (anonymous):

you split seven into 4 and 3

OpenStudy (nincompoop):

yeah we use that type of method in earlier days of factoring :P

OpenStudy (nincompoop):

HS algebra

OpenStudy (anonymous):

ok

OpenStudy (nincompoop):

can you try the rest?

OpenStudy (anonymous):

the problem is over

OpenStudy (anonymous):

i mean finished

OpenStudy (nincompoop):

not quite, we still have the absolute value, and we also have the higher square root

OpenStudy (anonymous):

2-root 3

OpenStudy (nincompoop):

okay!

ganeshie8 (ganeshie8):

isn't it redundant to put absolute bars as the square root already implies that you need to take positive value ?

OpenStudy (nincompoop):

yeah, it is

OpenStudy (anonymous):

well, yeah , but that's what is the questioon

ganeshie8 (ganeshie8):

\[\large \sqrt{(x-3)^2} = \sqrt{(3-x)^2} = |x-3| = |3-x|\]

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

yeah but this notion wasn't even introduced to me when I was taking algebra in HS. I learned this from reading Spivak's Calculus

OpenStudy (anonymous):

that is a fundamental idea

OpenStudy (nincompoop):